Project

General

Profile

2018-03-06-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

6th March 2018; 15:30 GMT
RAL, R1 2nd Floor East, MICE Meeting room

Phone details

Present: AK, CW, KL, ABr,

  1. Introduction: KL
    • This meeting: revised paper planning; and disscussion of decommissioning.
    • Had intended to make a test ramp of the magnets today. This was not possible because Hall was not ready.
  2. Notes of previous meeting and status of actions: All
    • Notes accepted. Status of actions:
      • CW: Organise test of QD/QP systems for SSs.
        • Ongoing. Tracker removal now complete. PRY completion schedule ramp for next week. CW will circulate the schedule.
      • CW: Bracing bars. CW to discuss with JB.
        • Done. To be installed but the studing etc. has been agreed.
      • CR: Will make list of magnet settings for field mapping exercise.
        • Done. A second iteration on the currents will be done by the end of the week.
      1. Hall status, stability of systems, planning for decommissioning and making safe to be complete by 31Mar18: CW
        • New information is that much of the space in R9 will be lost next week. Need to complete prepration for this.
        • Tracker cryostats can be disassembled and scrapped to recover the cryocoolers.
        • TOF, EMR, KL all off the beam line. FC off the beamline. Space now an issue and equipment needs to be removed and stored or scrapped.
      2. Status of preparations for magnetic survey measurements: CW
        • CERN mapping team will be here to start work on 19Mar18. Will need to run magnets during week of 19Mar18.
        • Preparation of magnet systems next week; CW will issue schedule. Presently believe Tuesday 13Mar18 will be test ramp.
        • Will discuss stepping between currents; identified need to minimise ramping time.
      3. Requests for equipment repatriation: CW
        • Now have equipment requests from TOF, KL, EMR, cryos for magnets to US, some other equipment requires to be sent.
        • CW: Will draft short document to define repatriation.
      4. Software and computing status: DR
        • No report this week.
      5. Paper status and plans: CR
        • See table uploaded: 2018-03-05-Papers-in-progress-v1.pdf . Points noted:
          • Emittance paper: systematic uncertainties and data MC comparison are remaining issues.
          • LiH scattering: Asymmetry in scattering distribution principle issue.
          • Emittance evolution: Working through details of systematics.
          • Emittance exchange: Now need to study beam selection and limiting appertures.
          • Other papers: see table; clear goals for completion of many papers. Papers on field-on scattering etc. need to progress before completion dates can be defined.
      6. AOB
        • None.

Summary of actions:

  • CW: Organise test of QD/QP systems for SSs.
  • CW: Will draft short document to define repatriation.

Back to MEMO:

Back

Back to MIPO:

Back

Documents

2018-03-05-Papers-in-progress-v1.pdf (37.3 KB) Long, Kenneth, 06 March 2018 16:00